OpenStudy (lilsis76):

Phy 101 A rectangular plate has a length of (21.3 +/- .2 ) cm and a width of (9.8 +/- .1) cm. Calculate the aea of the plate, including the uncertainty.

OpenStudy (aaronq):

\(A=l*w=21.3~cm*9.8 ~cm=208.74~cm^2\) \(\dfrac{s_A}{A}=\sqrt{(\dfrac{s_l}{l})^2+(\dfrac{s_w}{w})^2}\) \(\dfrac{s_A}{208.74~cm^2}=\sqrt{(\dfrac{0.2}{21.3})^2+(\dfrac{0.1}{9.8})^2}=2.894\) \(s_A=2.894~cm^2\)

OpenStudy (lilsis76):

lol Hi @aaronq could you explain what the s^A means, and what you did to get the answer. I want to be sure to understand. THe book is confusing me

OpenStudy (aaronq):

\(\huge s_x\) is the uncertainty in measurement x. so \(\huge s_A\) is the uncertainty in the area. Likewise, \(\huge s_l\) and \(\huge s_w\) are the uncertainties given for the length and the width, respectively. Looked where i subbed in the values, after that it's just algebra.

OpenStudy (lilsis76):

oh...okay

OpenStudy (aaronq):

you know, i just noticed i wrote something incorrectly. ignore the 2.984 in the third line, it shouldn't be there.

OpenStudy (lilsis76):

okay, im still studying what you did exactly. let me see if I have it right.

OpenStudy (lilsis76):

OKAY @aaronq i got SA= 2.893 thank you so much. UGH! math haha

OpenStudy (aaronq):

awesome!